# Equivariant concordance of periodic 2-knots in $S^{4}$

### v1

@@ 3.1. Definition and related notions
    $\varphi(K):=\frac{\sigma(X)-D\cdot D}{8}\mod 2$
    where $D\cdot D$ here represents the normal Euler number of $D$ relative to the Seifert framing of $K$.
@@
    We say $K$ and $J$ are band pass equivalent if there exists a sequence of planar isotopies, Reidemeister moves, and band passes taking a diagram of $K$ to a diagram of $J$. Band pass equivalence is an equivalence relation on the set of knots in $S^{3}$, and divides knots into two classes: those band pass equivalent to the unknot, and those band pass equivalent to the trefoil. We see therefore that the band pass equivalence class is an invariant of $K$.
Theorem 3.2 (Robertello).
Most recent treatments of the Arf invariant of knots use the definition (1); for instance, this is the definition given in [Sav12], Chapter 9. However, there are also several extensions and related invariants in 3- and 4-manifold topology that build on these equivalent definitions. For instance, the Arf invariant was successfully extended to knots in integer homology 3-spheres, where it interacts with the Rokhlin invariant of the ambient 3-manifold, see [Gor75]. There is also a notion of the Arf invariant of a properly embedded characteristic surface in a 4-manifold. A modern discussion of this invariant can be found in [Sav12]. However, most comprehensive account of all of these invariants and their relation to each other is given by Klug [Klu21], where he proves the following for a characteristic surface $F$ properly embedded in a smooth 4-manifold $X$ with $\partial F=K$ and $\partial X$ an integer homology 3-sphere:
$\operatorname{Arf}(F)+\operatorname{Arf}(K)\equiv\frac{\sigma(X)-F\cdot F}{8}+\mu(\partial X)\mod 2.$
Here $\mu$ denotes the Rokhlin invariant.
Below we give a generalization of the Arf invariant to null-homologous knots in arbitrary oriented 3-manifolds based primarily on definitions (2) and (3). Throughout, all knots are null-homologous in their ambient 3-manifolds and therefore have well-defined Seifert longitudes. For any embedded surface $F$ in a 4-manifold bounded by such a knot, we denote by $F\cdot F$ the normal Euler number of $F$ relative to the Seifert framing on the boundary.
Definition 3.3.
Let $K$ be a closed, null-homologous knot embedded in a 3-manifold $Y$. Let $X$ be a spin cobordism between $Y$ and $S^{3}$. Let $C$ be a properly embedded characteristic annulus in $X$ with normal Euler number $C\cdot C\equiv 0$ mod 16 such that $\partial C=K\sqcup J$, where $J$ is a knot in $S^{3}$. We define the Arf invariant of $K$ in $Y$, denoted $\operatorname{Arf}(K,Y)$, to be:
$\operatorname{Arf}(K,Y):=\operatorname{Arf}(J)\in\mathbb{Z}/2\mathbb{Z}$
To aid in our discussion of this invariant, we will call an annulus satisfying the conditions in Definition 3.3 an Arf annulus.
3.2. Well-definedness
To show that $\operatorname{Arf}(K,Y)$ is well-defined, it is helpful to establish the existence of spin fillings with Arf annuli for all null-homologous knots in 3-manifolds.
@@ 3.2. Well-definedness
For all null-homologous $K\subset Y$, there exists a spin cobordism $X$ between $Y$ and $S^{3}$ and an Arf annulus $C$ bounded on one end by $K$.
@@
By a result of Kaplan [Kap79], for any spin structure on $Y$ there always exists a spin cobordism between $Y$ and $S^{3}$ consisting of only 2-handles, and so we may interpret this cobordism as the result of spin surgery on a collection of framed curves in $Y$. Since $K$ is null-homologous, we may assume that $K$ is also null-homologous in the complement of the surgery curves, and in particular that $K$ bounds a surface $F$ in $Y$ disjoint from all the surgery curves. (To ensure this, we can start with any surface bounded by $F$ and introduce more genus to remove intersections with the surgery curves.) Thus we can obtain $C$ by tracing $K$ through successive cross-sections of the surgery: if $X$ is given by $Y\times I\cup\{2\text{-handles}\}$, then $C$ is given by $K\times I$. Furthermore, since none of the surgery curves intersect $F$, the product $F\times I$ embeds properly in $X$, and so the Seifert framings on $K$ and $J$ extend over $C$. We conclude that $C$ is 0-framed rel. boundary. ∎
@@
Figure 7. Schematic of an Arf annulus.
We also state a corollary of Theorem 3.2 that applies only to spin manifolds.
Corollary 3.5.
Let $K$ be a knot in $S^{3}$. Let $X$ be a spin filling of $S^{3}$ and let $D$ be a characteristic disk in $X$ bounded by $K$. Then:
$\operatorname{Arf}(K)\equiv\frac{D\cdot D}{8}\mod 2.$
Proof.
Rokhlin’s theorem [Roh52] states that spin 4-manifolds have vanishing signature mod 16. Thus by the equivalence of definitions (2) and (3) in Definitions 3.1, we obtain the result. ∎
With this established, well-definedness follows almost immediately. In the following proposition, for a manifold or knot $M$, M-M indicates $M$ with the opposite orientation.
Proposition 3.6.
Let $C_{1},C_{2}$ be two Arf annuli in different spin cobordisms $X_{1},X_{2}$ from $Y$ to $S^{3}$ as in Proposition 3.4, and let $J_{1}$, $J_{2}$ be their other boundary components, respectively. Then $\operatorname{Arf}(J_{1})=\operatorname{Arf}(J_{2})$.
Proof.
Without loss of generality, assume for the purposes of contradiction that $\operatorname{Arf}(J_{1})=1$ and $\operatorname{Arf}(J_{2})=0$. Let $X$ be obtained by gluing $X_{1}$ and $X_{2}$ along $Y$, and let $C$ be the annulus obtained by gluing $C_{1}$ and $C_{2}$ along $K$. Then $X$ is a spin cobordism from $S^{3}$ to $S^{3}$, and $C$ is an Arf annulus from $J_{1}$ to $J_{2}$. Since $\operatorname{Arf}(J_{2})=0$, we have $\operatorname{Arf}(-J_{2})$ = 0 as well. Thus by Proposition 3.4 and Corollary 3.5, we can find a spin filling $X^{\prime}$ of the $X_{2}$-boundary component of $X$ in which $J_{2}$ bounds a characteristic disk $D$ with $D\cdot D\equiv 0$ mod 16. So $X^{\prime}$ is a spin filling of $S^{3}$ in which $C\cup D$ is a characteristic disk with self-intersection 0 mod 16 bounded by $J_{1}$. But $\operatorname{Arf}(J_{1})=1$, which contradicts Corollary 3.5. ∎
It is worth spelling out the following corollary to our above work, in the precise wording that we will eventually need in Section 5.
Corollary 3.7.
Let $K$ be a null-homologous knot in a closed 3-manifold $Y$. Suppose that $\operatorname{Arf}(K,Y)=1$. Then there does not exist any pair $(D,X)$ where $X$ is a spin 4-manifold with $\partial X=Y$ and $D$ is a properly embedded, characteristic disk in $X$ with $D\cdot D\equiv 0$ mod 16 such that $D=\partial K$.
3.3. Extension to arcs and additivity
Definition 3.8.
Let $A$ be an arc properly embedded in a 3-manifold $Y$ such that $\partial Y\cong S^{2}$ and $A$ is null-homologous in $Y$ rel. boundary. Consider the union of $(A,Y)$ with $(\ell,B^{3})$ along their common boundary, where $\ell$ denotes the unknotted arc in $B^{3}$. We define $\operatorname{Arf}(A,Y)$ to be $\operatorname{Arf}(A^{\prime},Y^{\prime})$, where $A^{\prime}=A\cup\ell$ and $Y^{\prime}=Y\cup B^{3}$.
Lemma 3.9.
Let $K_{1}$ and $K_{2}$ be null-homologous knots in 3-manifolds $Y_{1}$ and $Y_{2}$, respectively. Choose 3-balls $B_{i}$ embedded in $Y_{i}$ such that $B_{i}$ intersects $K_{i}$ in a trivial tangle, and form the connected sum of pairs $(K,Y)=(K_{1}\#K_{2},Y_{1}\#Y_{2})$ along the boundary of $Y_{i}\setminus B_{i}$. Then:
$\operatorname{Arf}(K,Y)=\operatorname{Arf}(K_{1},Y_{1})+\operatorname{Arf}(K_{2},Y_{2})\mod 2.$
Proof.
Let $X_{1}$ and $X_{2}$ be spin cobordisms of $Y_{1}$ and $Y_{2}$ to $S^{3}$, with embedded Arf annuli for $K_{1}$ and $K_{2}$ terminating in knots $J_{1}$ and $J_{2}$, respectively. Then so long as the balls $B_{i}$ are sufficiently small, we may remove their trace through the cobordisms $X_{i}$ to obtain spin cobordisms $X_{i}^{\prime}$ between $Y_{i}\setminus B_{i}$ and $S^{3}\setminus B^{3}$, with embedded copies of $I\times I$ cobounded by the tangles of $K_{i}$ and $J_{i}$ together with the trace of the intersections of $K_{i}$ with $\partial B_{i}$. We then glue these two cobordisms with boundaries together to form a 4-manifold $X$ with boundary $Y\sqcup S^{3}$, with an annulus $C$ cobounded by $K$ in $Y$ and $J:=J_{1}\#J_{2}$ in $S^{3}$. It is then simple to check that $C$ is an Arf annulus, and so the result follows by the additivity of the classical Arf invariant. ∎
The above proof follows exactly the same for null-homologous arcs in 3-manifolds with spherical boundary.
4. Equivariantly slice 2-knots
In Section 2, we constructed for each periodic 2-knot a corresponding periodic Seifert 3-manifold $Y$ such that $\operatorname{Fix}(\tau)\cap Y=A$, with $A$ a single arc. Together with the definitions established in the previous section, we may define our invariant for periodic 2-knots.
Definition 4.1.
Let $S$ be a $d$-periodic 2-knot in $S^{4}$ such that $S$ bounds an equivariant Seifert 3-manifold $Y$ with fixed set $A$. Let $\overline{Y}$ be the quotient of $Y$ by $\tau$, and let $\overline{A}$ be the image of $A$ under the quotient map. The quantity $\alpha$ of $(S,Y,A)$ is then defined by:
$\alpha(S,Y,A):=\operatorname{Arf}(\overline{A},\overline{Y})\in\mathbb{Z}/2\mathbb{Z}$
We have not proven at this point that $\alpha$ is actually an invariant of $S$. The proof of this fact will follow from work in Section 5. However, we will be able to show in this section the following theorem, which states that periodic knots for which $\alpha$ vanishes for some Seifert 3-manifold are indeed slice.
Theorem 4.2.
Let $S$ be a $d$-periodic 2-knot, and let $Y$ be an equivariant Seifert 3-manifold with fixed arc $A$. If $\alpha(S,Y,A)=0$, then $S$ is equivariantly slice, regardless of the order $d$ of $\tau$.
The strategy to constructing the slice ball is as follows: First, we quotient $Y$ by $\tau$ to obtain some other 3-manifold $\overline{Y}$, as in the definition of $\alpha$. This realizes $Y$ as the cyclic cover of $\overline{Y}$ branched over an arc $\overline{A}$, where $A$ descends to $\overline{A}$ under the branched covering map. We then isotope $\overline{Y}$ into the interior of $\overline{B^{5}}$ and ambiently surger the pair $(\overline{Y},\overline{A})$ to $(\overline{B^{3}},\overline{\ell})$ where $\overline{\ell}$ is an unknotted arc. In order to ensure that these surgeries can be performed ambiently without introducing new intersections with $\operatorname{Branch}(\tau)$, we will require $\alpha$ to vanish. We then lift the isotopy and surgery curves and disks to $Y\subset B^{5}$, obtaining a sequence of suitable equivariant surgeries taking $(Y,A)$ to $(B^{3},\ell)$, since the latter is its own cyclic branched cover.
We now proceed with the proof of Theorem 4.2, which will occupy the next few subsections. The first step of this procedure is to take the quotient of $Y$ by $\tau$. Note that by construction, $Y\cap\operatorname{Fix}(\tau)=A$, where $A$ is a single arc. Thus the quotient $Z:=Y/(y\sim\tau(y))$ is a topological 3-manifold $\overline{Y}$. In particular, if $\pi$ is the quotient map $Y\to Z$, then the $n$-fold cyclic cover of $\overline{Y}$ branched along $B:=\pi(A)$ recovers $Y$.
The following lemma allows us to push $\overline{Y}$ into $\overline{B^{5}}$, and by extension $Y$ into $B^{5}$, to prepare it for surgery. We will not actually use the extra properties of equivariant Morse position and the handle structure of the fixed set in this section, but it will be important later in Theorem 5.3. Note that in order to talk about equivariant isotopy of $Y$, we must realize it as the image of some equivariant map $\psi\colon M\to B^{5}$, where $M$ is an abstract 3-manifold of the same homeomorphism type as $Y$. Note $M$ then comes equipped with an action $\rho$ of $\mathbb{Z}/d\mathbb{Z}$ given by the pullback of $\tau$ by $\psi$.
Lemma 4.3.
For a fixed periodic 2-knot $S$, let $\psi\colon M\to B^{5}$ be a proper embedding whose image is the Seifert 3-manifold $Y$ as constructed in Proposition 2.1. Then $\psi$ is equivariantly isotopic rel. boundary to an equivariant proper embedding $\psi^{\prime}$ such that there exists an equivariant Morse function $h$ on $B^{5}$ with respect to which $\psi^{\prime}(M)$ is in ordered equivariant Morse position and $h|_{A}$ has only one critical point, where $A$ denotes $\psi^{\prime}(M)\cap\operatorname{Fix}(\tau)$.
Proof.
Let $\overline{M}$ be the quotient of $M$ by $\tau$, and let $\overline{\psi}$ be the induced map to $\overline{B^{5}}$ with image $\overline{Y}$. Choose a relative Morse function $g$ on $\overline{M}$ (which is by definition identically 0 on $\partial M$) which is valued in $[0,1]$ achieves its maximum on an interior point of $\overline{\psi}^{-1}(\overline{A})$.
Fix a local equivariant Morse function $\overline{h}$ on a neighborhood $U$ of $\overline{Y}$ in $\overline{B^{5}}$ (including some of the interior of $B^{5}$, not just on the boundary) such that $h|_{\overline{\partial B^{5}}}\equiv 0$. We can then construct an isotopy $\overline{\Psi}\colon\overline{M}\times[0,1]\to h^{-1}([0,1])$ given by:
$\overline{\Psi}(x,t)=(\overline{\psi}(x),t\cdot g\circ\overline{\psi}(x))$
Then $\overline{\Psi}(x,0)=\overline{\psi}(x)$ in the bounding $\overline{S^{4}}$, and $\overline{\Psi}(x,1)$ is a proper embedding $\overline{Y}$ of $\overline{M}$ in $\overline{B^{5}}$ such that $\overline{\Psi}(\overline{A})$ has exactly one maximum. Here again we use Proposition 1.11 to rearrange $\overline{Y}$ via ambient isotopy to ordered equivariant Morse position without changing the single critical point of $\overline{A}$.
We may then take the $d$-fold branched cover of $\overline{B^{5}}$ and $\overline{Y^{\prime}}$ inside it to get $Y$ pushed into $B^{5}$ in ordered equivariant Morse position with respect to the lifted equivariant Morse function $h$ on $B^{5}$. Since $\overline{A}$ had only a single maximum with respect to $\overline{h}$, $A$ has only a single maximum with respect to $h$. As before, this isotopy lifts by Schwarz’s Isotopy Lifting Theorem [Sch80]. ∎
As the reader will see later in the following section, it is crucial that the above lemma allows us to preserve the single critical point of $\overline{A}$.
4.1. Spin structures and ambient surgery in dimension 5

### v2

@@ 5.1. The concordance invariant for annuli
    if $\alpha(N)=1$, then $N$ is concordant to the one-twist spun trefoil minus its spin axis, and
@@
    every Montesinos annulus in $S^{1}\times B^{3}$ satisfies exactly one of the above two conditions.
Thus the smooth concordance group of Montesinos annuli $\mathcal{C}_{2}^{\circ}$ is isomorphic to $\mathbb{Z}/2\mathbb{Z}$.
We start by showing that $\alpha$ is a well-defined quantity independent of choice of $Y$ and $A$, for which we need the following technical lemma.
Lemma 5.3.
Let $S$ be a 2-knot in $S^{4}$ and let $R$ be a smoothly unknotted 2-sphere in $S^{4}$ such that $S\cap R$ consists of two points. Let $Q$ be a smoothly unknotted 3-sphere embedded in $S^{5}$ such that $Q\cap S^{4}=R$. Let $Y$ be a Seifert 3-manifold for $S$ such that $Y\cap R$ consists of a null-homologous arc $A$, and consider $Y$ as the image of an embedding $\psi\colon M^{3}\to S^{4}$ for some 3-manifold with spherical boundary $M$. Then $\psi$ is isotopic rel. boundary to a proper embedding $\psi^{\prime}$ of $M$ in $B^{5}$ such that there exists a Morse function $f$ on $B^{5}$ with respect to which $\psi^{\prime}(M)$ is in ordered Morse position. Furthermore, the arc $A^{\prime}\colon=\psi^{\prime}(M)\cap Q$ can be taken to be fully vertical with respect to $f|_{\psi^{\prime}(M)}$ outside of a single critical point.
Proof.
Choose a relative Morse function $g:M\to[0,1]$ such that:
  1. (1)
    $g|_{\partial M}\equiv 0$,
  2. (2)
    $\psi(g^{-1}(1))$ is a single point in $A$, and
  3. (3)
    away from $g^{-1}(1)$, $\psi^{-1}(A)$ is fully vertical with respect to $g$.
Fix a local Morse function $f$ on a half-ball $U$ containing $Y$ in $B^{5}$} such that $f|_{\partial B^{5}}\equiv 0$ and $Q$ is fully vertical with respect to $f$. (Parametrizing $B^{5}$ and $Q$ as in Proposition 4.2 will do.) We can then construct an isotopy $\Psi\colon M\times[0,1]\to f^{-1}([0,1])$ given by:
$\Psi(x,t)=(\psi(x),t\cdot g\circ\psi(x))$
Then $\Psi(x,0)=\psi(x)$ in the bounding $S^{4}$}, and $\Psi(x,1)$ is a proper embedding $Y^{\prime}$ of $M$ in $B^{5}$} such that $f|_{A^{\prime}}$ for $A^{\prime}:=Y^{\prime}\cap Q$ has exactly one maximum and is otherwise fully vertical with respect to $f|_{Y^{\prime}}$ by our choice of $g$ and $\Psi$. Here again we use Proposition 3.5 to rearrange $Y$ via ambient isotopy to ordered Morse position without changing the single critical point of $f|_{A^{\prime}}$ or the full verticality of $A^{\prime}$ with respect to $f$ away from this critical point. ∎
This lemma shows that it is possible to perturb Seifert 3-manifolds into the 5-ball in such a way that their intersections with the removed ball $Q$ is fully vertical with respect to the restriction of a Morse function $f$ to $Y$. This will enable us to use the constructions in Propositions 4.2 and 4.3 to show that the $\alpha$ invariant is well-defined.
@@
The value of $\alpha(N)$ is independent of the choice of $Y$ and $A$.
@@
Let $S$ be the corresponding sphere for $N$. Let $Y_{1}$ and $Y_{2}$ be two Seifert 3-manifolds for $S$, each intersecting the removed sphere $R$ in $S^{4}$ in null-homologous arcs $A_{1}$ and $A_{2}$. By the abuse of notation, in this proof we will speak of isotoping $Y_{i}$ and $A_{i}$ rel. boundary. $A_{i}$ always refers to the intersection of the 3-manifold $Y_{i}$ with $Q$, and is always topologically an arc.
Consider the $S^{4}$ in which $S$ is embedded as the equator of a copy of $S^{5}$, such that this $S^{4}$ bounds two copies $B_{1}$ and $B_{2}$ of $B^{5}$ on either side. In this setup, we view $R$ as the equator of a smoothly unknotted 3-sphere $Q$ in $S^{5}$.
Parametrize a 5-dimensional neighborhood $U$ of $S$ with coordinates $(x,y,z,w,t)$ so that the $S^{4}$ containing $S$ intersects $U$ in the subspace defined by setting $t=0$. Then by Lemma 5.3, for $i=1,2$ we may isotope $Y_{i}$ rel. boundary into $B_{i}$ such that each each arc of intersection $A_{i}$ of $Y_{i}$ with $Q$ is fully vertical with respect to the function $f\colon U\to\mathbb{R}$ given by projection onto the $t$-axis when restricted to $Y_{i}$. By Proposition 3.5, may then isotope each $Y_{i}$ rel. boundary to ordered Morse position with respect to $f$ such that this isotopy is supported away from $A_{i}$, maintaining the full verticality of the $A_{i}$’s with respect to $f|_{Y_{i}}$.
Let $Y$ be the union of $Y_{1}$ and $Y_{2}$ along $S$, and $K$ be the union of $A_{1}$ and $A_{2}$ along their boundary. Then $Y$ is a closed 3-manifold in $S^{5}$ which intersects $Q$ in a knot $K$ which is fully vertical with respect to $f|_{Y}$ except for exactly two critical points and is null-homologous in $Y$, since each of the $A_{i}$’s was null-homologous in its $Y_{i}$. Thus $Y$, AA, and $Q$ satisfy the conditions of Proposition 4.3, and hence $\operatorname{Arf}(K,Y,\mathfrak{s}_{0})=0$ where $\mathfrak{s}_{0}$ is the restriction of the ambient spin structure on $S^{5}$ to $Y$. But $(K,Y,\mathfrak{s}_{0})$ is the connected sum of pairs $(A_{1},Y_{1},\mathfrak{s}_{0})\#(A_{2},Y_{2},\mathfrak{s}_{0})$, and so by Lemma 2.14, we have that $\operatorname{Arf}(A_{1},Y_{1},\mathfrak{s}_{0})=\operatorname{Arf}(A_{2},Y_{2},\mathfrak{s}_{0})$. ∎
We similarly deduce the following proposition, which is one direction of the classification of Montesinos annuli up to concordance. The idea is to use the same trick of building a 3-manifold intersecting a 3-sphere in a fully vertical knot as in the previous proposition, but this time, one of the two 3-manifolds glued together is a slice ball for $S$ which is not necessarily boundary parallel, and thus can’t on its own be used to compute $\alpha$.
Proposition 5.5.
Let $N$ be an annulus in $S^{1}\times B^{3}$ and suppose that $N$ is slice. Then $\alpha(N)=0$.
Proof.
Parametrize a neighborhood of $N$ in $S^{5}$ with coordinates as in the previous proposition. Let $f$ be a local Morse function on this neighborhood which agrees with projection onto the $t$-coordinate.
Since $N$ is slice, there exists a smooth embedding $\psi$ of $S^{1}\times I\times I$ into $S^{1}\times B^{3}\times I$ such that $\psi(-,0)=N$, $\psi(-,1)=N_{U}$ the unknotted Montesinos annulus, and $\psi$ is level-preserving on $S^{1}\times S^{0}\times I\subset S^{1}\times I\times I$. We may therefore embed $S^{1}\times B^{3}\times I$ into $S^{5}$ such that the final $I$ factor agrees with the $t$-coordinate in our parametrization. Under this embedding, the vertical boundary $S^{1}\times S^{0}\times I$ of the concordance is fully vertical with respect to $f$. Thus the union $C$ of the corresponding spheres for the annuli in each $t$-slice intersect the union of the removed spheres in each $t$-slice in two arcs $A_{+}$ and $A_{-}$ which are fully vertical with respect to $f|_{C}$. Let $Q$ be an unknotted 3-sphere in $S^{5}$ such that its intersection with our parametrized neighborhood is the $(z,w,t)$-subspace, and note that $Q\cap C$ is precisely $A_{+}\cup A_{-}$.
Now since $N_{U}$ is unknotted, we may find an unknotted 3-ball $B$ bounded by its corresponding sphere $S_{U}$ in the $t=1$ slice which intersects $Q$ in an unknotted arc. By Lemma 5.3, we may find an isotopy of $B$ rel. boundary to some 3-ball $B^{\prime}\subset S^{5}$ such that $B^{\prime}\cap Q$ is an unknotted arc which is fully vertical with respect to $f|_{B^{\prime}}$.
Now let $D$ be the 3-ball in $S^{5}$ given by the union $B^{\prime}\cup C$. Then $\partial D$ is the corresponding sphere $S$ for $N$ embedded in the $t=0$ slice. Note that since $A_{+},A_{-}$, and $B^{\prime}\cap Q$ were all fully vertical with respect to $f$, $D\cap Q$ is a single arc $\ell$ which is unknotted in $D$ and is fully vertical with respect to $f|_{D}$.
Finally, let $Y$ be a Seifert 3-manifold for $S$ such that $Y$ intersects $Q$ in a null-homologous arc $A$, and isotope $Y$ into the $t\leq 0$ half of $S^{5}$ by Lemma 5.3 such that its intersection with $Q$ is fully vertical with respect to $f|_{Y}$. The union $Y\cup D$ is then a closed 3-manifold embedded in $S^{5}$ such that its intersection $A\cup\ell$ with $Q$ is a null-homologous knot which is fully vertical with respect to $f|_{Y\cup D}$ apart from two critical points. We may then apply Proposition 4.3 to conclude $\operatorname{Arf}(A\cup\ell,Y\cup D,\mathfrak{s}_{0})=0$, and hence by the additivity of the Arf invariant established in Lemma 2.14, we have that $\operatorname{Arf}(A,Y,\mathfrak{s}_{0})=\operatorname{Arf}(\ell,D,\mathfrak{s}_{0})$. But since $D$ is a 3-ball and $\ell$ is unknotted, $\operatorname{Arf}(\ell,D,\mathfrak{s}_{0})=0$, and so $\alpha(N):=\operatorname{Arf}(A,Y,\mathfrak{s}_{0})=0$ as well. ∎
Finally, we establish the additivity of the $\alpha$ invariant, which follows quickly from the additivity of the Arf invariant.
Lemma 5.6.
Let $N$ and $N^{\prime}$ be Montesinos annuli. Then $\alpha(N\natural N^{\prime})=\alpha(N)+\alpha(N^{\prime})$.
Proof.
Let $Y$ and $Y^{\prime}$ be Seifert 3-manifolds for the corresponding spheres $S$ and $S^{\prime}$ and let $A$ and $A^{\prime}$ be null-homologous arcs of intersection of $Y,Y^{\prime}$ with removed spheres $R,R^{\prime}$. Then by the stacking construction given in Definition 1.9, we see that the partial boundary connected sum $Y$ of $Y$ and $Y^{\prime}$ forms a Seifert 3-manifold for $N\natural N^{\prime}$ which intersects the connected sum of the removed spheres in the boundary connected sum of the tangles $A$ and $A^{\prime}$. The result then follows from Lemma 2.14. ∎
5.2. Constructing slice balls
The final step in the classification of Montesinos annuli up to concordance is to actually build the concordances. We do this by building concordances of the corresponding spheres while continuously keeping track of their intersection with the removed 3-ball $Q$. We will work in the same parametrized neighborhood as used in the constructions in Section 4.
The general strategy of the construction is the same as that of Kervaire and Sunukjian: build a Seifert manifold $Y$ for $S$, push it into the interior of the 5-ball, and then ambiently surger it inside of $B^{5}$ to a 3-ball. However, because we want our final slice ball $B$ for $S$ to intersect $Q$ in an unknotted arc $\ell$ so that we may remove $Q$ to obtain a slice cylinder for $N$, we must keep track of the intersection of $Y$ with $Q$ at each stage of the surgery. This imposes an extra restriction that differs from the usual obstruction to ambient surgery in dimension 5, discussed in Subsubsection 5.2.2.
@@ 5.2. Constructing slice balls
Figure 12. A schematic of the setup in preparation for surgery. Example surgery curves are shown.
5.2.1. Spin structures and ambient surgery in dimension 5